# SRPO: A Cross-Domain Implementation of Large-Scale Reinforcement Learning on LLM

### v1

@@ Abstract
Recent advances of reasoning models, exemplified by OpenAI’s o1 and DeepSeek’s R1, highlight the significant potential of Reinforcement Learning (RL) to enhance the reasoning capabilities of Large Language Models (LLMs). However, replicating these advancements across diverse domains remains challenging due to limited methodological transparency. In this work, we present two-Staged history-Resampling Policy Optimization (SRPO), which successfully surpasses the performance of DeepSeek-R1-Zero-32B on the AIME24 and LiveCodeBench benchmarks. SRPO achieves this using the same base model as DeepSeek (i.e. Qwen2.5-32B) and relies solely on RL, without prior Supervised Fine-Tuning (SFT). Building upon Group Relative Policy Optimization (GRPO), we introduce two key methodological innovations: (1) a two-stage cross-domain training paradigm designed to balance the development of mathematical reasoning and coding proficiency, and (2) History Resampling (HR), a technique to address ineffective samples. Our comprehensive experiments validate the effectiveness of our approach, dedicating to offer valuable insights into scaling LLM reasoning capabilities across diverse tasks.
Figure 1: Benchmark performance of SRPO, achieving superior results with only 20% of DeepSeek’s training steps. The values shown are pass@1 scores, averaged over 32 samples per question.
@@ 1 Introduction
Combining the aforementioned techniques, SRPO achieves 50.0 pass@1 on AIME24 and 41.6 pass@1 on LiveCodeBench [14], surpassing the performance of DeepSeek-R1-Zero-Qwen-32B [3] (47.0 points on AIME & 40.2 on LiveCodeBench), within 2,000 steps — just 1/5 of the R1-Zero training steps. In the latter part of the article, we provide a detailed analysis of the reasoning patterns in large-scale RL training on math and coding tasks.
@@ 2.1 GRPO
The core idea of GRPO [13] is to estimate the baseline through a relative reward within a group of rollouts. Therefore, it reduces the computational cost of the critic model and improves the training stability. Specifically, for each question q, the model generates a group of responses $o_{1},o_{2},...,o_{G}$ and calculates the corresponding rewards $r_{1},r_{2},...,r_{G}$. $A_{i}$ is the advantage obtained by normalizing the rewards within each group.
@@ 3.1 Challenges with Vanilla GRPO Implementation
    Intrinsic Response-Length Conflict between Math and Code: We find that math problems often elicit longer and meticulous step-by-step reasoning patterns (long CoT), while RL training on code data does not encourage the model to generate longer responses. Directly mixing these two types of data creates an intrinsic conflict, limits the model’s ability to optimally develop either the necessary verbosity required for mathematical reasoning or the succinct conventions preferred in coding tasks. Consequently, this conflict results in suboptimal performance in both domains.
@@ 3.2 Two Stage Training
    Stage 2 (Skill Integration): In this stage, the coding data is introduced into the training process. Using the reasoning foundation established in Stage 1, this stage was aimed at gaining proficiency in code generation. Beyond directly enhancing the coding capabilities, this stage also aims to elicit procedural thinking, recursion, and tool-calling capabilities from the model.
@@ 3.2.1 Comparative Analysis of Training Strategies
Training only on math data led to a consistent increase in response length and strong performance on math benchmarks. Importantly, this fosters robust reasoning abilities that generalize well; when presented with coding tasks, the model attempts detailed step-by-step reasoning. Observed behaviors include meticulous step checking and re-examination within mathematical problem-solving processes. This reflects the characteristic of mathematical data of eliciting reasoning abilities.
@@ 3.3 History Resampling (HR)
  • Retain ”Informative”: Samples that yield either mixed outcomes (both correct and incorrect rollouts) or exclusively incorrect outcomes are retained. These samples exhibit positive reward variance, ensuring nonzero advantages and effective gradient signals. Additionally, for hard samples where all rollouts were incorrect within the current epoch, we keep them in the dataset. The rationale is that some of these initially challenging problems may become relatively easier for the updated policy, thereby generating positive advantages in subsequent rollouts. This underlying idea of this strategy aligns with curriculum learning[16], progressively exposing the model to samples that are, on average, more challenging, to improve training efficiency.
@@
Compared to dynamic sampling method[12], which adjusts the rollout number on the fly, history sampling significantly improves computational efficiency. The positive impact of history resampling on the growth of response length is illustrated in Figure 9.
@@ 4.4 Thinking Behavior Analysis
We are surprised to find that in the course of RL training, the frequency of self-reflection, correction, and backtracking of the model gradually increases. It demonstrates that the model shows the "self-verification" ability. We believe that the emergence of "reflection" analogous to human cognitive processes represents an adaptive behavior exhibited by the model during policy optimization in RL.

### v2

@@ Abstract
00footnotetext: Model available at: https://huggingface.co/Kwaipilot/SRPO-Qwen-32B/
Recent advances of reasoning models, exemplified by OpenAI’s o1 and DeepSeek’s R1, highlight the significant potential of Reinforcement Learning (RL) to enhance the reasoning capabilities of Large Language Models (LLMs). However, replicating these advancements across diverse domains remains challenging due to limited methodological transparency. In this work, we present two-Staged history-Resampling Policy Optimization (SRPO), which surpasses the performance of DeepSeek-R1-Zero-32B on the AIME24 and LiveCodeBench benchmarks. SRPO achieves this using the same base model as DeepSeek (i.e. Qwen2.5-32B), using only about 1/10 of the training steps required by DeepSeek-R1-Zero-32B, demonstrating superior efficiency. Building upon Group Relative Policy Optimization (GRPO), we introduce two key methodological innovations: (1) a two-stage cross-domain training paradigm designed to balance the development of mathematical reasoning and coding proficiency, and (2) History Resampling (HR), a technique to address ineffective samples. Our comprehensive experiments validate the effectiveness of our approach, offering valuable insights into scaling LLM reasoning capabilities across diverse tasks.
Figure 1: SRPO achieves superior results with only 10% of DeepSeek’s training steps. The values shown are pass@1 scores, averaged over 32 samples per question.
@@ 1 Introduction
Combining the aforementioned techniques, SRPO achieves 50.0 pass@1 on AIME24 and 41.6 pass@1 on LiveCodeBench [14], surpassing the performance of DeepSeek-R1-Zero-Qwen-32B [3] (47.0 points on AIME & 40.2 on LiveCodeBench), training for over 1K steps — just about 1/10 of the R1-Zero training steps. In the latter part of the article, we provide a detailed analysis of the reasoning patterns in large-scale RL training on math and coding tasks.
@@ 2.1 GRPO
The core idea of GRPO [13] is to estimate the baseline through a relative reward within a group of rollouts. This approach reduces the computational cost of the critic model and improves the training stability. Specifically, for each question q, the model generates a group of responses $o_{1},o_{2},...,o_{G}$ and calculates the corresponding rewards $r_{1},r_{2},...,r_{G}$. $A_{i}$ is the advantage obtained by normalizing the rewards within each group.
@@ 3.1 Challenges with Vanilla GRPO Implementation
    Intrinsic Response-Length Conflict between Math and Code: We find that math problems often elicit longer and meticulous step-by-step reasoning patterns (long CoT), while RL training on code data does not encourage the model to generate longer responses. Directly mixing these two types of data creates an intrinsic conflict and limits the model’s ability to optimally develop either the necessary verbosity required for mathematical reasoning or the succinct conventions preferred in coding tasks. Consequently, this conflict results in suboptimal performance in both domains.
@@ 3.2 Two Stage Training
    Stage 2 (Skill Integration): In this stage, the coding data is introduced into the training process. Using the reasoning foundation established in Stage 1, this stage aims at gaining proficiency in code generation. Beyond directly enhancing the coding capabilities, this stage also aims to elicit procedural thinking, recursion, and tool-calling capabilities from the model.
@@ 3.2.1 Comparative Analysis of Training Strategies
Training only on math data led to a consistent increase in response length and strong performance on math benchmarks. Importantly, this fosters robust reasoning abilities that generalize well; when presented with coding tasks, the model attempts detailed step-by-step reasoning. Observed behaviors include meticulous step checking and re-examination within mathematical problem-solving processes. This reflects the ability of mathematical data to elicit reasoning behaviors.
@@ 3.3 History Resampling (HR)
  • Retain ”Informative”: Samples that yield either mixed outcomes (both correct and incorrect rollouts) or exclusively incorrect outcomes are retained. These samples exhibit positive reward variance, ensuring nonzero advantages and effective gradient signals. Additionally, for hard samples where all rollouts were incorrect within the current epoch, we keep them in the dataset. The rationale is that some of these initially challenging problems may become relatively easier for the updated policy, thereby generating positive advantages in subsequent rollouts. The underlying idea of this strategy aligns with curriculum learning[16], progressively exposing the model to samples that are, on average, more challenging, to improve training efficiency.
@@
Compared to the dynamic sampling method[12], which adjusts the rollout number on the fly, history sampling significantly improves computational efficiency. The positive impact of history resampling on the growth of response length is illustrated in Figure 9.
@@ 4.4 Thinking Behavior Analysis
We are surprised to find that in the course of RL training, the frequency of self-reflection, correction, and backtracking of the model gradually increases. This demonstrates that the model develops "self-verification" ability. We believe that the emergence of "reflection" analogous to human cognitive processes represents an adaptive behavior exhibited by the model during policy optimization in RL.